OpenStudy (anonymous):

cos(3pi/4+pi/6)= (-root6 - root2)/4 right?

OpenStudy (kagıtucak):

=cos135.cos30-sin135.sin30=(-root2/2).root3/2-(root2/2).1/2=-root6/4-root2/4= (-root6-root2)/4 you are right

OpenStudy (anonymous):

thanks